

********************************************************************
UARS WEEKLY STATUS REPORT - 13 Sep 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
On Saturday, 9/5/02, SUSIM resumed solar measurements after the turn
off for the yaw-around. An extra day was needed to warm up
sufficiently. Afterward, SUSIM has successfully performed its daily
mid-resolution and MgII solar scans during periods of predicted UARS
telemetry. High Resolution (0.15 nm) scans have resumed; we expect
to do these about twice per month.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 31 August 2002 (Orbit 59993, GMT Day 243) through
06 September 2002 (Orbit 60098, GMT Day 249).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM(Z/H) 245/1938 60035 POWER OFF
SUSIM 246/2123 60036 POWER OFF
SOLSTICE 245/2229 60037 TO DAY ONLY MODE
HRDI 246/0008 60038 LOW POWER MODE
SOLSTICE 249/1409 60092 POWER ON - DAY ONLY MODE
SUSIM 239/1548 60093 POWER ON
A Reverse-to-Forward Yaw Around maneuver (number 112) was successfully
executed in Orbit 60062 beginning 247/15:03:42Z and completing at
247/15:40:52Z on 04 September 2002. All instruments were in their off
or standby mode (including HRDI and PEM). Due to the SA position at
269 degrees, the effective night length of the spacecraft after the
maneuver was approximately 56.52 minutes. Post maneuver analysis of
data indicate that during the Yaw Around, the end-of-night load bus
voltage was 25.76 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 20.51 and 20.04 percent
respectively. The next (forward-to-reverse) YAW is scheduled for 10
October 2002 (DOY 283).
The beta angle ranged between the angles of 16.2 degrees to 7.7
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 57 degrees on 23 September 2002 (DOY 266).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 0.95 degrees C. Additional
battery performance data for each battery on 30 August and 06
September 2002 are provided in Appendix D.
The clock error ranged from -10.1 to 10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
PEM = 01
SUSIM = 02
SOLSTICE = 02
HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Raytheon development team has delivered a new software delivery,
which takes care of several minor problems that exist on the current
Eclipse system. This patch has yet to be tested on the prototype
system due to a hard disk failure, but should be tested within the
next couple weeks.
MEETINGS
The FOT attended the UARS Monthly Status Meeting on 05 September 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 October 2002 (DOY 283).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 06 September 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
60018 244/1613 02-306 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
60035 245/1938 02-307 PEM(Z/H) POWER OFF
60036 245/2123 02-308 SUSIM POWER OFF
60036 245/2123 02-310 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
60037 245/2229 02-309 SOLSTICE TO DAY ONLY MODE
60038 246/0008 02-311 HRDI LOW POWER MODE
60061 247/1345 02-312 MPS VT TO LEVEL-6
60062 247/1441 02-312 OBC/TAM TAM TO MON MODE/YAW MEM MONITOR
60062 247/1520 02-312 OBC/TAM SAFEHOLD DIR CHG/TAM TO UPDATE
MEMORY MONITOR TO NORMAL OPS
60076 248/1329 02-312 MPS VT TO LEVEL-5
60079 248/1736 02-313 MPS BATSOC RESET
60092 249/1409 02-314 SOLSTICE POWER ON - DAY ONLY
60093 249/1548 02-315 SUSIM POWER ON (PART I)
______________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 06 September 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 06 September 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60018 244/1613 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
60036 245/2123 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 06 September 2002
30 August 2002 - GMT Day 242 Beta = 20.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -134.4/-156.8 -3.05/-4.62 +00.0/+00.0 08.5 0.000 0.0
2 +11.2/-44.8 +5.47/+3.55 +28.0/-13.6 24.2 1.040 19.1
3 +22.4/-39.2 +4.51/+2.60 +26.8/-13.2 24.2 1.030 18.4
06 September 2002 - GMT Day 249 Beta = 07.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -134.4/-156.8 -3.68/-4.94 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-100.8 +4.51/+3.87 +28.4/-12.8 25.4 1.020 19.5
3 +22.4/-22.4 +3.87/+2.92 +27.2/-12.8 25.4 1.010 19.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
60061 247/1345 L5 -> L6
60076 248/1329 L6 -> L5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 06 September 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-046 FOT 59831 232/0534 30824 Data dropout due to an
equipment fail-over at WSGT.
02-047 SSPP 59844 233/0213 N/A SSPP in wrong mode
02-048 FOT N/A 242/1300 N/A Loss of UPS-CMS connectivity
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-045 FOT 59717 224/1458 14170 Generic Late Acquisition